A Large, cylindrical birthday cake has a diameter of 50cm. it is cut into twelve equal slices. what is the area of icing on top of each slice, to nearest square centimetre?
1
Expert's answer
2012-09-18T12:11:29-0400
The area of the top of the cake is

A = π(D/2)² = π(50/2)² ≈ 1963.4954 cm².

So, the top area of the slice is

S = A/12 ≈ 1963.4954/12 = 163.6246 cm² ≈ 164 cm².